Q1.a) Find the direction and magnitude of an electric field that exerts a westward force on an electron.b) what magnitude and direction force does this field exert on a proton
Q2. An electron gun has parallel plates separated by 4.00cm and gives electrons 25.0KeV of energy
a) what is the electric field strength between the plates. b)what force will this field exert on a piece of plastic with a 0.500μF charge that gets between the plates?
Q3. Find the total capacitance of the combination of capacitors. C1=1μF C2 =5μF C3=8 μF
1
Expert's answer
2020-02-24T10:46:52-0500

1)

a) As you have not mentioned the magnitude of force exerted on the electron so lets us assume that force exerted on the electron is equal to F so the electric field will be equal to E=F/e where e is the charge on one electron

as the electron carries a negative charge so the force experienced by electron will in opposite direction of the electric field so if the electric field exerts a westward force on the electron then electric field must be in the opposite direction i.e. in the eastward direction

b) As the proton carries positive charge and of magnitude same as electron so it will experience force F in eastward direction


2) Voltage = 25*103 eV/e =25*103 V

Electric field = voltage/distance = 25*103/0.04 = 62500 V/m


3)net capacitance if they are in parallel"=C_1+C_2+C_3=(1+5+8)\mu F=14\mu F"